Fermi National Laboratory


Accelerator Update

Friday December 19
The day shift began with Operations monitoring store 3103, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 10:32 AM, Tevatron (TeV) experts began moving the new collimators at sector A48 into the beam.

Recycler experts successfully sent five shots of antiprotons to the Main Injector (MI) from 12:11 PM to 12:41 PM. Pbar resumed stacking.

At 1:18 PM, a Switchyard (SY) expert began a study period.

Operators noticed that the current single-store record for integrated luminosity, set back on July 25, was being broken. At 2 PM it was at 2145 nb-1 and still climbing. It eventually reached 2244 nb-1.

Operations terminated store 3103 at 4 PM. Operators made D-zero (D0) ready for access. Cryo system technicians began work on three wet engines around the TeV.

Cryo techs completed their wet engine work by 5 PM and D0 came out of access at 5:40 PM.

Operations put the TeV through a dry squeeze at 5:54 PM.

At 6:30 PM, Operation established a 12x0 study-store and them began manipulating the A48 collimators.

Operations put the TeV into shot setup at 8 PM.

At 8:04 PM, a Linac quadrupole power supply (QPS511) tripped off. Operators tried to reset the device but could not keep it up and running. Operators called in two Linac techs to investigate a possible ground fault.

Saturday December 20
The midnight shift began with beam off due to a ground fault investigation in Linac, with the TeV in a dry squeeze, and with the Recycler circulating beam.

At 1 AM, the Linac techs reported that L:OPS511 was possibly ground faulted. By 1:22 AM a Linac expert had tuned beam with another unused quad to compensate for the bad quad.

Operators put the TeV back into shot setup at 1:55 AM.

Operations began sending beam to MiniBooNE at 2:35 AM.

Operations established store 3108 at 4:19 AM. Its initial luminosity was 51.79E30; this set another new luminosity record.

At 6:15 AM, the TeV suffered a six-house quench. The store was lost.

Operators discovered, at 6:40 AM, that the power was out at the TeV sector A0 Kicker building. A TeV expert said that a power loss at A0 would have made all the kickers fire, which would have caused the quenches.

At 7:30 AM, operators made D-zero (D0) ready for an access.

Operators made the TeV transfer hall ready for access at 8:52 AM. Techs will investigate the abort kicker problem.

By 12:58 PM electricians had found and disconnected an unused welding receptacle that had caused the breaker trip. A TeV expert started conditioning the A0 kickers.

Vacuum techs added several vacuum pumps to the TeV sector B area at 1:46 PM to help the slow vacuum recovery.

At 6:30 PM, vacuum techs discovered that the TeV sector B14 cell was saturated with helium, which was why the vacuum wasn't recovering. Operations began warming the cryo system for sector B1.

Vacuum techs intentionally spoiled the B1 vacuum at 9:52 PM.

Sunday December 21
The midnight shift began with TeV sector B14 warming up and with the Recycler storing antiprotons.

At 1:38 AM, the Cryo Coordinator said the B1 house would take about forty hours to reach room temperature. (Later this estimate was raised to seventy hours since the low betas also had to be warmed.)

Operations put the Recycler into shot setup at 1 PM and eventually made nineteen transfers of antiprotons.

Pbar resumed stacking at 4:32 PM.

Monday December 22
The midnight shift began with TeV warm-up underway, with Pbar conducting reverse proton studies, with Switchyard studies, with Recycler storing antiprotons, with D0 in access, and with MiniBooNE taking beam.

At 5:26 AM, Pbar completed their studies and resumed stacking.

Operators turned off beam for MiniBooNE and readied Booster for an access at 7 AM.

Today's Plans
The plan for today and for the next twelve days, or so, is to repair the TeV vacuum leak in sector B. Hopefully, our next store will be early January of next year, 2004.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/22/2003   email Fermilab